# Back-Reaction of Super-Hubble Fluctuations, Late Time Tracking and Recent Observational Results

### v1

@@ III Methods and Data Analysis
where the “fid” denotes the quantities measured in the fiducial cosmology, which is the FRW model used to convert redshifts into distances in the survey. In the above, $r_{d}$ is the comoving sound horizon evaluated at the redshift at which baryon-drag optical depth is equal to unity, and it depends on quantities at redshifts higher than that. In the case of the back-reaction model presented in Section II, the quantity $r_{d}$ will be considered to be equal to the fiducial one, since we are here not investigating the effects of back-reaction at earlier times. Therefore in our analysis we are going to consider the approximate expression [^8] For a similar analysis in a different context, see for instance the work of RefClifton:2024mdy ,
@@
where $\Delta H\equiv H(z)-H^{fid}(z)$. We are going to use information from the DESI first-year galaxy and quasar results DESI:2024uvr , and the Lyman-$\alpha$ forest results DESI:2024lzq to interpret the constraints on $\alpha_{||}$ in terms of constraints on $\Delta H$ at given redshifts.
@@
We first solve Eq.9 in order to obtain H(z) predicted by the back-reaction model. We then plot the resulting function $\Delta H/H_{\Lambda CDM}$ together with the observational values and error bars for $\Delta H/H_{\Lambda CDM}$ shown in Table I. However, in order to plot the quantity $\Delta H/H_{\Lambda CDM}$ predicted by the back-reaction model, the parameters in equation 7 must be fixed. As mentioned earlier, the phase is fixed in such a way that $\rho_{BR}$ is zero at redshift $z=3.400$. The amplitude $A$ and the ”frequency”, $nz$, remain to be fixed. We will focus on values of the amplitude which are of the order of $A\sim 10^{-1}-10^{-2}$. Larger values of the amplitude would give rise to too large oscillations of the Hubble parameter, while smaller values would not lead to observable effects. We cannot exactly determine the amplitude from the considerations in Section II since this would require a non-perturbative calculation that has not yet been achieved. On the other hand, the considerations in Section II imply that the frequency of oscillation is set by a fraction of the Hubble rate. Hence, we would expect that a non fine-tuned frequency would correspond to a value of $n$ with order of magnitude approximately equal to one.
We show the results obtained in the next Section.
IV Results
@@
In this section we show the predictions of the back-reaction model compared with the data of DESI:2024lzq ; DESI:2024uvr . In Fig. 1 we show the late-time behavior of the function $\Delta H/H_{\Lambda CDM}$ (curves in color) for values of the parameter $A$ ranging from zero to $A=-0.05$ and for a fixed value of $n\approx 5$ in Eq.9, compared with the DESI results (red dots and error bars). Different shapes of the theoretical curve can be obtained using different values of the frequency and amplitude. We choose, just for the sake of illustration, particular sets of values that provide a shape for the curves which tracks the DESI data better than the standard $\Lambda$CDM model. Note that the lowest redshift bin is in disagreement with both the $\Lambda$CDM model and our scenario. At the present time the error bars on the data points are still too large to draw any definite conclusions, but our results show that it might be possible for this model to accommodate future improved data (assuming that the central values of the data points remain the same but the error bars shrink). Note that the values of the free parameters (amplitude and frequency) which we have used are within the range of values that can be considered as “natural”. In particular, the frequency is in the range expected from qualitative arguments.
@@
Figure 1: $\Delta H/H$ predicted by the back-reaction model compared with DESI data
Although the results look promising, our preliminary analysis should not be viewed as conclusive in terms of statistical evidence. With the next data releases the error bars are expected to shrink, and then the data should be able to provide us with a more solid conclusion on the preference of the data for $\Lambda$CDM or a dynamical dark energy. The objective of Fig. 1 is to show the range of amplitudes of the oscillating dark energy term which can be in agreement with the DESI constraints. We can see that, for the chosen value $n\sim 5$, values up to $|A|\sim 0.05$ can fit the data within the observational uncertainties. As mentioned earlier, the only data that our model, as well as $\Lambda$CDM, does not fit, corresponds to the lowest redshift point, whose validity is still being discussed.
As a preliminary analysis, we performed a test minimizing the $\chi^{2}$ of the model with respect to the DESI data, using several different minimization methods and testing several combinations of initial values. The method consists of executing a set of minimization routines, with each minimization sequence having, as initial guesses for the parameters $A$ and $n$, the values of a pair that is one of the possible combinations between previously chosen values of $A$ and $n$. We chose the initial values of $A$ to lie within the range $[-0.025,1]$, and those of $n$ within $[1,4]$ These intervals correspond to initial guesses of suitable parameter values and are starting points for the parameter exploration. In principle one could also vary the phase of the oscillations. However, as mentioned previously, the phase was chosen such that the model is equivalent to the standard model when $z=3400$, which is approximately the beginning of the matter-dominated phase. This is justified by the fact that the oscillatory density of the effective component arising from back-reaction is taken to be proportional to the energy density of matter, and therefore another parametrization should be chosen before the onset of the matter era. We are letting the back-reaction effects build up beginning with the value $\rho_{BR}=0$ at $z\sim 3400$. For considerations of late time cosmology this should be fine.
The expression to be minimized is the $\chi^{2}$ of $\Delta H(z)/H_{\Lambda CDM}$ (which expresses the relative difference between the proposed model and the standard cosmological model as a function of redshift) with respect to the experimental data from DESI. As described previously, the DESI experimental results were interpreted in terms of $\Delta H$, and the error propagation was calculated. Some minimization results lead to a shape more in line with the one indicated by DESI data than others. However, in practically all the cases the $\chi^{2}$ obtained from $\sim 500$ minimizations performed were smaller than the one of the $\Lambda$CDM model . On the other hand, this analysis were restricted to amplitudes with order of magnitude $A=\mathcal{O}(10^{-2})$ or smaller. For large values of $A$ the fit is poor. This is shown in Fig.2, where the value of $\chi^{2}$ is plotted as a function of the amplitude. The value of $n$ chosen for each point in Fig. 2 was the one that provided the best fit for the respective value of the amplitude. From this same Figure we can also see that for higher values of the amplitude the back-reaction model provides a fit worse than the standard model. We can also see from Fig. 2 that the minimum value of the $\chi^{2}$ is shifted with respect to the $\Lambda$CDM value (which is recovered for the value $A=0$). While the error bars still do not allow us to conclude that our model performs better than $\Lambda$CDM, these results can be viewed as an indication of the importance of further studies of this model, rather than a firm statistic conclusion.
@@
Figure 2: The value of $\chi^{2}$ for the back-reaction model with respect to DESI data as a function of the parameter $A$.
In Figs 3 and 4 we plot, respectively, the deceleration parameter and the equations of state of the back-reaction model and of the total content of the model, according to the equations in Section II. The pair of values $(A\sim-0.03,n\sim 4.997)$ was chosen as an illustrative example in all these figures.
In Fig.3 we show the behavior of the deceleration parameter predicted by the back-reaction model for the same values of the parameters $A$ and $n$. We can see that this model is able to correctly predict the transition from decelerated to accelerated expansion around the expected redshift.
Figure 3: Deceleration parameter from $z=2.5$ to $z=0$ predicted by the back-reaction model.
In Fig.4 (left) we illustrate the evolution of the equation of state for the effective back-reaction fluid, given by Eq.13. We considered again the same values for the parameters $A$ and $n$. We can see in the plot at the left that the equation of state shows the average value $w\sim 0$. One can also see that, at the exact redshifts when $\rho_{BR}$ crosses zero, a spike can be seen in the plot of $w_{BR}$. This should be expected from the fluid equation of state $w_{BR}\equiv p_{BR}/\rho_{BR}$. The spikes correspond to the points where the quantity $\Delta H/H_{\Lambda CDM}$ crosses zero in Fig.1. However, this behavior does not correspond to a problem  [^9] A similar behavior can be seen for instance, in the model analyzed in Ref.Tiwari:2024gzo . since the observable quantities in cosmology correspond to integrals of the equation of state over a period of time. Therefore these spikes are not expected to be present in any cosmological observable. In addition, the contribution from the back-reaction effective fluid is always subdominant compared to the other components in the Universe. The effective equation of state of the total fluid has the expected overall behavior, as shown in Fig.4 (right), i.e., it oscillates around the $\Lambda CDM$ prediction.
Another important aspect to be analyzed is whether the back-reaction model considered here can provide some relief to the Hubble tension, and if so, for which values of the parameters. In order to shed some light to this question we perform an analysis of the value of $H_{0}$ predicted for different combinations of the parameters $n$ and $A$. The analysis uses the method described previously. We show in Fig. 5 a plot of $H_{0}$ as a function of $A$, $H_{0}=H_{0}(A,n_{\text{best}})$, where $n_{\text{best}}$ is the value of $n$ chosen from the minimization method performed. As in the previous analysis, we fix our model to be equal to $\Lambda CDM$ at $z=3.400$. We can see in the plot that, for the representative chosen values of the parameters, the $H_{0}$ predicted by our model is very close to the $\Lambda$CDM one. This is partly due to the choice of the fixed A and n and also due to the fact that the back-reaction contribution decreases with $\rho_{m}$, causing the difference with respect to $\Lambda$CDM to decrease at later times.
Figure 5: $H_{0}$ as a function of $A$, for $n=n_{\text{best}}$ chosen from the minimization method performed.
In order to see what could be the highest value of $H_{0}$ that this model could predict, we can investigate the Hubble constant value for an amplitude as high as $A=1$. From all the results of our minimizations, the highest value that was obtained was $H_{0}=1.14\;H_{0\Lambda CDM}$, which is achieved for $A=1$ and $n=2$. For this same value $n=2$, but now for a smaller amplitude, $A=0.5$, we found values for $H_{0}$ as high as $H_{0}=1.075\;H_{0\Lambda CDM}$, which is already more than enough to solve the tension problem. We can see that it can be possible to solve the Hubble tension in our back-reaction model for sufficiently high values of the amplitude. However, from Fig 1 we can see that such high values of $A$ does not provide a good fit to DESI data. For the amplitude $A=\pm 0.05$, which is approximately the highest value that fits DESI error bars for the chosen frequency, we obtain for the Hubble constant of the back-reaction model $H_{0}/H_{0\Lambda CDM}=1\pm 10^{-3}$, i.e., it’s basically indistinguishable from $\Lambda CDM$.
It’s important to emphasize again that a priori there is no reason that the amplitude $A$ should be independent of time. In fact, from the considerations of back-reaction it is not unreasonable to expect that the effective amplitude of the oscillating contribution to DE performs damped oscillations. In this case, $A$ could be large at early times and then slowly decay to a value consistent with the current late time data. This extended scenario might therefore be able to provide a solution of the Hubble tension. On the other hand, without a careful analysis of the fluctuations produced in the model, it is not possible to gauge whether the scenario would be consistent with other constraints, e.g. the value of $\sigma_{8}$.
V Conclusion
In the present work, we studied a toy model of oscillating dark energy motivated by considerations of the back-reaction of super-Hubble fluctuations in the late universe. We proposed a parametrization of the oscillating dark energy component to account for the back-reaction effect on the background cosmology (based on the conjectured dynamical scaling fixed point solution describing the back-reaction effect at a fully non-perturbative level). We computed the Hubble parameter at low redshifts predicted by the model, obtaining an oscillatory dynamics. We compared the resulting background cosmology with data from the radial component of the Baryon Acoustic Oscillations obtained from the recent DESI survey Jiang:2024xnu . We found that, for a frequency scale expected from theoretical considerations, a fit to the recent observations can be obtained for a certain range of the amplitude of the oscillation, which is a free parameter in the analysis, and for a frequency of oscillation motivated by the theory. At the present time the observational error bars are still so large that there is no statistically significant advantage of our model over the standard $\Lambda$CDM cosmology. But, with upcoming observations the error bars will shrink and may lead to a way to extract a non-vanishing amplitude of an oscillating component to Dark Energy in a reliable way.
We also investigated the impact of our model on the Hubble tension. We found that for a large amplitude $A$ of the oscillatory component (such that the energy density in the oscillating term is comparable to the matter energy density) a sufficiently large change in $H_{0}$ results to explain the Hubble tension. However, if the amplitude $A$ is constant in time, such a large value of $A$ is inconsistent with the low redshift data. If, on the other hand, $A$ depended on time and decreased from a large value near the time of recombination to a small value at the present time, the Hubble constant tension might be successfully adressable. However, detailed studies of the effects of our model on cosmological fluctuations is required to check the consistency of the scenario.
The next step should in fact be to analyze the perturbations in our scenario. This would allow us to make comparisons with other observable quantities, e.g. CMB anisotropies, and in particular the Integrated Sachs Wolf (ISW) effect. In this context, in Ref.Pace:2011kb the behavior of several cosmological observables, such as the linear growth factor, the ISW effect, the number counts of massive structures, and the matter and cosmic shear power spectra were estimated for several oscillating dark energy models. It was shown that, for several of the models considered, then independently of the amplitude and the frequency of the dark energy oscillations, none of these observables showed a significant oscillating behavior as a function of redshift. This is a consequence of the fact that such observables are integrals over functions of the expansion rate along the cosmic history, consequently smoothing oscillatory features below the level of detectability. This analysis was performed under certain assumptions, for instance, supposing that Dark Energy does not cluster on the scales of interest. In addition, the models studied in this reference have a different behavior than the back-reaction model we consider in our work. A complete analysis specific for our scenario is hence still missing, but we believe it might be possible that the main conclusions of the work of Pace:2011kb might also hold for our model. As the observables mentioned are integrals over functions of the expansion rate along the cosmic history, we expect that also in our scenario the oscillatory features will be smoothed, keeping the ISW effect in agreement with current observations.
Our work can be considered to be a first-step analysis limited to the background predictions of the model. Nevertheless, in order to obtain more conclusive results, a study of the predictions of the model for linear perturbations is required. Our results, although preliminary, highlights the importance of considering the back-reaction effect of super-Hubble fluctuations in the late Universe, as such back-reaction effects are expected to be present in any cosmological scenario with entropy fluctuations. We have shown that this provides a dynamics which is interesting in light of recent observational results.

### v2

@@ III Methods and Data Analysis
where the “fid” denotes the quantities measured in the fiducial cosmology, which is the Planck-normalized $\Lambda$CDM model used to convert redshifts into distances in the survey  [^11] Using the notation of Ref.DESI:2025zgx , we can write $\alpha_{||}=\alpha_{\text{iso}}\alpha_{\text{AP}}^{2/3}$.. In the above, $r_{d}$ is the comoving sound horizon evaluated at the redshift at which baryon-drag optical depth is equal to unity, and it depends on quantities at redshifts higher than that. In the case of the back-reaction model presented in Section II, the quantity $r_{d}$ will be considered to be equal to the fiducial one, since we are here not investigating the effects of back-reaction at earlier times. Therefore in our analysis we are going to consider the approximate expression [^12] For a similar analysis in a different context, see for instance the work of RefClifton:2024mdy ,
@@
where $\Delta H\equiv H(z)-H^{fid}(z)$. We are going to use information from the DESI DR2 galaxy, quasar and Lyman-$\alpha$ results DESI:2025zgx to interpret the constraints on $\alpha_{||}$ in terms of constraints on $\Delta H$ at given redshifts.
@@
Tracer Redshift $\alpha_{||}$ $\Delta H/H^{fid}$
LRG1 $0.51$ $0.9615419512\pm 0.01922057331$ $0.03999622561\pm 0.0207888212$
LRG2 $0.706$ $0.9647661889\pm 0.01642269702$ $0.03652057001\pm 0.01764413334$
ELG2 $1.321$ $0,9932025202\pm 0.01256687344$ $0.006844001718\pm 0,01273947749$
QSO $1.484$ $1.02675147\pm 0.02099900114$ $-0.02605447469\pm 0.01991902012$
Lya $2.33$ $1.008212343\pm 0.01700856044$ $-0.008145449517\pm 0.01673260419$
LRG3 $0.922$ $0.9953917692\pm 0.04052811979$ $0.004629564906\pm 0.04090424355$
ELG1 $0.955$ $1.001814039\pm 0.01652471677$ $-0,001810754569\pm 0,01646492654$
Table 1: Dataset used for $\alpha_{||}$ from the DESI DR2 galaxy, quasar and Lyman-$\alpha$ results taken from Ref.DESI:2025zgx , and its interpretation in terms of $\Delta H/H^{fid}$ at given redshifts.
Our procedure is the following: We first solve Eq.10 (with $\Lambda_{0}$ taken to be the same as in the fiducial cosmology) in order to obtain H(z) predicted by the back-reaction model. From this we can plot the resulting function $\Delta H/H_{\Lambda CDM}$ for the model together with the observational values and error bars shown in Tab.1. However, in order to plot the quantity $\Delta H/H_{\Lambda CDM}$ predicted by the back-reaction model, the parameters in equation 7 must be fixed. As mentioned earlier, the phase is fixed in such a way that $\rho_{BR}$ is zero at redshift $z=3.400$. The amplitude $A$ and the ”frequency”, $nz$, remain to be fixed. We will focus on values of the amplitude which are of the order of $A\sim 10^{-1}$. Larger values of the amplitude would give rise to too large oscillations of the Hubble parameter, while smaller values would not lead to observable differences compared to our fiducial cosmology. We cannot exactly determine the amplitude from the considerations in Section II since this would require a non-perturbative calculation that has not yet been achieved. On the other hand, the considerations in Section II imply that the frequency of oscillation is set by a fraction of the Hubble rate. Hence, we would expect that a non fine-tuned frequency would correspond to a value of $n$ with order of magnitude approximately equal to one.
We show the results obtained in the next Section.
IV Results
In this section we show the predictions of the back-reaction model compared with the data of Ref.DESI:2025zgx . In Fig. 1 we show the late-time behavior of the quantity $\Delta H/H_{\Lambda\text{CDM}}$ (curves in color) as predicted by the back-reaction model, for values of the amplitude ranging from zero (which represents $\Lambda\text{CDM}$) to $A=-0.15$ and for a fixed frequency characterized by the parameter $n\approx 2.4$ introduced in Eq.10. These predictions are compared with the DESI DR2 results (red dots and error bars). Different shapes of the theoretical curve can be obtained using different values of the frequency and amplitude. We choose, just for the sake of illustration, examples of values that provide a shape for the curves which tracks the DESI data better than the standard $\Lambda$CDM model. In addition, in choosing the frequency value to be illustrated in the Figures, we selected models which can alleviate the Hubble tension even slightly, and we disregard models with too high frequencies ($n\gtrsim 5$) since they could lead to smaller $\chi^{2}$ simply due to overfitting. Among the frequency values that satisfy these criteria, we chose the one that provides the smaller $\chi^{2}$, which corresponds to $n\approx 2.4$. However, higher values for the frequency can be considered, also providing a $\chi^{2}$ value lower than the $\Lambda$CDM. Fig.6 in the Appendix shows the same plot as Fig.1 but for DR1 data. Comparing Fig.1 with Fig.6 we see that the evidence of an oscillating dark energy component is stronger in DR2 than in DR1.
Note that the values of the free parameters (amplitude and frequency) which we have used are within the range of values that can be considered as “reasonable” from the theoretical point of view. In particular, the frequency is in the range expected from qualitative arguments.
Figure 1: $\Delta H/H$ predicted by the back-reaction model compared with DESI DR2 data
Although the results look promising, our preliminary analysis should not be viewed as conclusive in terms of statistical evidence. With future data releases, the error bars are expected to shrink, and by combining those with datasets from other sources it should then be possible to obtain more solid conclusion on the preference of the data for $\Lambda$CDM or a dynamical dark energy. The objective of Fig. 1 is to show the range of amplitudes of the oscillating dark energy term which can be in agreement with the DESI constraints. We can see that, for the chosen value $n\sim 2.4$, values as low as $|A|\sim 0.05$ can fit the data within the observational uncertainties.
As a preliminary analysis, we performed a test minimizing the $\chi^{2}$ of the model with respect to the DESI DR2 data. The method used consists of executing a set of minimization routines, such that for each minimization sequence we chose the initial values of $A$ to lie within the range $[-0.1,0.1]$, and those of $n$ within $[3,5]$ These intervals correspond to initial guesses of suitable parameter values and are starting points for the parameter exploration.
The expression to be minimized is the $\chi^{2}$ of $\Delta H(z)/H_{\Lambda CDM}$ (which expresses the relative difference between the proposed model and the standard cosmological model as a function of redshift) with respect to the experimental data from DESI DR2. As described previously, the DESI experimental results were interpreted in terms of $\Delta H$, and the error propagation was calculated. Some minimization results lead to a shape more in line with the one indicated by DESI data than others. However, in practically all the cases considered, for the mentioned initial values, the $\chi^{2}$ obtained were smaller than the one of the $\Lambda$CDM model. On the other hand, these analysis were restricted to amplitudes with order of magnitude $A=\mathcal{O}(10^{-1})$ or smaller. For larger values of $A$ the fit is poor. This is shown in Fig.2, where the value of $\chi^{2}$ is plotted as a function of the amplitude. From this same figure we can also see that for values of the amplitude higher than zero and lower than $A\sim-0.15$ the back-reaction model provides a fit worse than the standard model (which is recovered for the value $A=0$, the star in the plot). We can also see from Fig. 2 that the minimum value of the $\chi^{2}$ is shifted with respect to the $\Lambda$CDM value. A similar plot using the data from DR1 can be seen in the Appendix. By comparing the $\chi^{2}$ as a function of the amplitude for the two datasets, DR1 and DR2, we see that the minimum value of the $\chi^{2}$ does not change considerably. However from the overall analysis, we obtain that the frequencies that tend to provided best fitting for the DR2 data are in general smaller then the ones that fits better the DR1 data. This general tendency seems to be in agreement with the results of Ref. Ormondroyd:2025iaf . While the error bars still do not allow us to conclude that our model performs better than $\Lambda$CDM, these results can be viewed as an indication of the importance of further studies of this model, rather than a firm statistical conclusion.
Figure 2: The value of $\chi^{2}$ for the back-reaction model with respect to DESI data as a function of the parameter $A$. The value for $\Lambda$CDM is represented by a star.
In Figs 3 and 4 we plot, respectively, the deceleration parameter and the equations of state (EOS) of the back-reaction model (the EOS for the effective backreaction fluid and the total EOS including all the components), according to the equations in Section II. The pair of values $(A\sim-0.08,n\sim 2.4)$ was chosen as an illustrative example in all these figures. We can see in Fig.3 that this model is able to correctly predict the transition from decelerated to accelerated expansion around the expected redshift.
Figure 3: Deceleration parameter from $z=2.5$ to $z=0$ predicted by the back-reaction model, together with the $\Lambda$CDM prediction.
In Fig.4 (left) we illustrate the evolution of the equation of state for the effective back-reaction fluid, given by Eq.14. We considered again the same values for the parameters $A$ and $n$. We can see in this plot that the equation of state shows the average value $w\sim 0$. One can also see that, at the exact redshifts when $\rho_{BR}$ crosses zero, a spike can be seen in the plot of $w_{BR}$. This should be expected from the fluid equation of state $w_{BR}\equiv p_{BR}/\rho_{BR}$. The spikes correspond to the points where the quantity $\Delta H/H_{\Lambda CDM}$ crosses zero in Fig.1. However, this behavior does not correspond to a problem  [^13] A similar behavior can be seen for instance, in the model analyzed in Ref.Tiwari:2024gzo . since the observable quantities in cosmology correspond to integrals of the equation of state over a period of time. Therefore these spikes are not expected to be present in any cosmological observable. In addition, the contribution from the back-reaction effective fluid is always subdominant compared to the other components of the Universe. The effective equation of state of the total fluid has the expected overall behavior, as shown in Fig.4 (right), i.e., it oscillates around the $\Lambda CDM$ prediction.
Another major challenge that has become more pressing following the DESI data is the Hubble tension problem Colgain:2025nzf ; Poulin:2025nfb ; Pang:2025lvh . Recent results from DR2 bolster the case for a dynamical dark energy scenario exhibiting phantom crossing behavior. This framework yields a lower inferred value of the Hubble constant, thereby intensifying the Hubble tension — which can be regarded as a symptom of potential underlying inconsistencies in the standard model. Therefore it is interesting to analyse in which cases, for which values of the parameters, the back-reaction model considered can provide some relief to the Hubble tension. In order to shed some light to this question we perform an analysis of the value of $H_{0}$ predicted for different combinations of the parameters $n$ and $A$. The analysis uses the method described previously. We show in Fig. 5 a plot of $H_{0}$ ($H_{0}=H_{0}(A,n_{\text{best}})$) as a function of $A$, where $n_{\text{best}}\sim 2.4$ is the value of $n$ chosen from the minimization method performed. We can see in the plot that, for the representative chosen values of the parameters, the $H_{0}$ predicted by our model is very close to the $\Lambda$CDM one. This is partly due to the choice of the fixed A and n and also due to the fact that the back-reaction contribution is proportional to $\rho_{m}$, therefore decreasing at later times, causing the difference of our model with respect to $\Lambda$CDM also to decrease.
Figure 5: $H_{0}$ as a function of $A$, for $n$ chosen from the minimization method performed.
From all the results of our minimizations, the highest value that was obtained was $H_{0}\approx 1.009\;H_{0\Lambda CDM}$, which is achieved for $A\approx-0.07$ and $n\approx 2.06$. For the amplitude $A=-0.05$, for instance, we obtain $H_{0}/H_{0\Lambda CDM}=1\pm 10^{-3}$, i.e., it’s basically indistinguishable from $\Lambda CDM$. Nevertheless, it could be possible to solve the Hubble tension given a sufficiently high absolute values of the amplitude. However, from Fig 1 we can see that such high absolute values of $A$ do not provide a good fit to DESI data. If we extrapolate and investigate the Hubble constant value for an amplitude as low as $A=-1$, we obtain $H_{0}\approx 1.09\;H_{0\Lambda CDM}$, considering $n=n_{\text{best}}$.
It is important to emphasize again that a priori there is no reason that the amplitude $A$ should be independent of time. In fact, from the considerations of back-reaction it is not unreasonable to expect that the effective amplitude of the oscillating contribution to DE performs damped oscillations. In this case, $A$ could be large at early times and then slowly decay to a value consistent with the current late time data. This extended scenario might therefore be able to provide a solution of the Hubble tension. On the other hand, without a careful analysis of the fluctuations produced in the model, it is not possible to gauge whether such scenario would be consistent with other constraints, e.g. the value of $\sigma_{8}$.
V Conclusion
In the present work, we studied a toy model of oscillating dark energy motivated by considerations of the back-reaction of super-Hubble fluctuations in the late universe. We proposed a parametrization of the oscillating dark energy component to account for the back-reaction effect on the background cosmology (based on the conjectured dynamical scaling fixed point solution describing the back-reaction effect at a fully non-perturbative level). We computed the Hubble parameter at low redshifts predicted by the model, obtaining an oscillatory dynamics. We also compared the resulting background cosmology with data from the radial component of the Baryon Acoustic Oscillations obtained from the recent DESI survey DESI:2025zgx . We found that, for a frequency scale expected from theoretical considerations, a fit to the recent observations can be obtained for a certain range of the amplitude of the oscillation, which is a free parameter in the analysis, and for a frequency of oscillation motivated by the theory. At the present time the observational error bars are still too large to observe a statistically significant advantage of our model over the standard $\Lambda$CDM cosmology. But, with upcoming observations the error bars will shrink and may lead to a way to extract a non-vanishing amplitude of an oscillating component to dark energy in a reliable way.
We also investigated the impact of our model on the Hubble tension. We found that for a large amplitude $A$ of the oscillatory component (such that the energy density in the oscillating term is comparable to the matter energy density) a sufficiently large change in $H_{0}$ results to explain the Hubble tension. However such a high amplitude does not provide a good fit to the data. If, on the other hand, $A$ depended on time and decreased from a large value near the time of recombination to a small value at the present time, the Hubble constant tension might be successfully addressable. However, detailed studies of the effects of our model on cosmological fluctuations is required to check the consistency of such a scenario.
The next step should in fact be to analyze the perturbations in our scenario. This would allow us to make comparisons with other observable quantities, e.g. CMB anisotropies, and in particular the Integrated Sachs Wolfe (ISW) effect. In this context, in Ref.Pace:2011kb the behavior of several cosmological observables, such as the linear growth factor, the ISW effect, the number counts of massive structures, and the matter and cosmic shear power spectra were estimated for several oscillating dark energy models. It was shown that, for several of the models considered, independently of the amplitude and the frequency of the dark energy oscillations, none of these observables showed a significant oscillating behavior as a function of redshift. This is a consequence of the fact that such observables are integrals over functions of the expansion rate along the cosmic history, consequently smoothing oscillatory features below the level of detectability. This analysis was performed under certain assumptions, for instance, supposing that dark energy does not cluster on the scales of interest. In addition, the models studied in that reference have a different behavior compared to the back-reaction model we consider in our work. A complete analysis specific for our scenario is hence still missing, but we believe it might be possible that the main conclusions of the work of Pace:2011kb might also hold for our model. As the observables mentioned are integrals over functions of the expansion rate along the cosmic history, we expect that also in our scenario the oscillatory features will be smoothed, keeping the ISW effect in agreement with current observations.
Our work can be considered to be a first-step analysis limited to the background predictions of the model. Our results, although preliminary, highlights the importance of considering the back-reaction effect of super-Hubble fluctuations in the late Universe, as such back-reaction effects are expected to be present in any realistic scenario. We have shown that this provides a dynamics which is interesting in light of recent observational results.
Appendix A Previous DESI DR1 Analysis
In this Appendix we show the analysis of the back-reaction model in light of the previous DR1 results for the sake of comparison. Using the DR1 data we produce the Table below with the values for $\alpha_{||}$ from Refs. DESI:2024uvr DESI:2024lzq , and its interpretation in terms of $\Delta H/H^{fid}$ at given redshifts.
@@ Appendix A Previous DESI DR1 Analysis
In Fig.6 we show the late-time behavior of the function $\Delta H/H_{\Lambda CDM}$ (curves in color) as predicted by the back-reaction model, for values of the parameter A ranging from zero to $A=-0.15$ and for a fixed value of $n\sim 5$. These predictions are compared with the DESI DR1 results (red dots and error bars). Different shapes of the theoretical curve can be obtained using different values of the frequency and amplitude. To illustrate the method, we select particular sets of parameter values that result in curve shapes that more closely align with the DESI DR1 data.
@@
Figure 6: $\Delta H/H$ predicted by the back-reaction model compared with DR1 data. In all the curves we considered the value $n\sim 5$.
We performed a test minimizing the $\chi^{2}$ of the model with respect to the DR1 data using the same method described in Section IV. In Fig.7, we show the value of $\chi^{2}$ plotted as a function of the amplitude. As in the analysis for DR2 shown in Fig.1, we can see that also in this case higher values of the amplitude the back-reaction model provides a fit worse than the standard model. Again we can see that the minimum value of the $\chi^{2}$ is shifted with respect to the $\Lambda$CDM value (which is recovered for the value A = 0). The minimum is achieved for an amplitude slightly smaller in modulus compared to the case with DR2 shown in Fig.2.
@@
Figure 7: The value of $\chi^{2}$ for the back-reaction model with respect to DR1 data as a function of the parameter $A$. The value for $\Lambda$CDM is represented by a star.
Although the best-fit frequency and amplitude of our model show slight variations when using DESI DR2 instead of DR1, and a modest improvement in the $\chi^{2}$ value is observed relative to the $\Lambda$CDM model when using DR2, the overall conclusions remain consistent across both DESI data releases.